OpenStudy (anonymous):

Indices, Surd and Numeral Systems #2 Question 42 A binary number has n digits and all digits are '1's. Convert the binary number into a decimal number in terms of n.

OpenStudy (amistre64):

lets try out a few small ones to see if theres a pattern

OpenStudy (amistre64):

1 = 1 = 2 - 1 11 = 3 = 4 - 1 111 = 7 = 8 - 1 the value appears to be 2^n -1

OpenStudy (anonymous):

Yes. But how to explain it?

OpenStudy (amistre64):

if you add 1 to it you get: 1000....n which is 2^n

OpenStudy (amistre64):

so subtract 1 from the 2^(n+1) spot

OpenStudy (anonymous):

Sorry... but I don't understand that. I understand the part ' add 1', but not the 'subtract one'.

OpenStudy (anonymous):

Why is it 2^(n+1), but not 2^n - 1?

OpenStudy (amistre64):

if you have n 1s to begin, that is one less in value than the a string of digits that starts with a 1 with n zeros after it

OpenStudy (amistre64):

maybe im mixing the n parts in my head ... tho

OpenStudy (amistre64):

2^0 = 0 2^1 = 10 2^2 = 100 2^3 = 1000 2^4 = 10000 so there are n zeros behind the 1 digit

OpenStudy (amistre64):

2^0 = 1 :)

OpenStudy (amistre64):

so if we subtract 1 from those we get: 2^0 -1 = 0 2^1 - 1 = 1 2^2 - 1 = 11 2^3 - 1 = 111 yeah, the (n+1) was a misthought

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

In this case, I should write when 111.....1 + 1, ^ n '1's It becomes, 100...0, that is 2^n ^ n '0's So, 111...1 = 100...0 - 1 = 2^n -1 ^n '1's ^n '0's Is that correct?